### v2

# supplementary

Abstract.
Epidemiological investigations of regionally aggregated spatial data often involve detecting spatial health disparities between neighboring regions on a map of disease mortality or incidence rates. Analyzing such data introduces spatial dependence among the health outcomes and seeks to report statistically significant spatial disparities by delineating boundaries that separate neighboring regions with widely disparate health outcomes. However, current statistical methods are often inadequate for appropriately defining what constitutes a spatial disparity and for constructing rankings of posterior probabilities that are robust under changes to such a definition. More specifically, non-parametric Bayesian approaches endow spatial effects with discrete probability distributions using Dirichlet processes, or generalizations thereof, and rely upon computationally intensive methods for inferring on weakly identified parameters. In this manuscript, we introduce a Bayesian linear regression framework to detect spatial health disparities. This enables us to exploit Bayesian conjugate posterior distributions in a more accessible manner and accelerate computation significantly over existing Bayesian non-parametric approaches. Simulation experiments conducted over a county map of the entire United States demonstrate the effectiveness of our method and we apply our method to a data set from the Institute of Health Metrics and Evaluation (IHME) on age-standardized US county-level estimates of mortality rates across tracheal, bronchus, and lung cancer.
1. Introduction
We propose methods for reporting spatial disparities on maps. Disparities in health outcomes is a rapidly evolving field within public health (see, e.g., rao2023, and references therein) and spatial variation in health outcomes play a crucial role in detecting health disparities. While the role of spatial data analysis in epidemiological investigations has been extensively documented (see, e.g., waller2004applied; waller2010handbook; lawson2013statistical; lawson2016handbook, and further references therein), inference pertaining to disparities comprises detecting “boundaries” on disease maps that identify significant differences among neighbors. These difference boundaries delineate regions with significantly different disease mortality and incidence rates and assists in improved geographic allocation of health care resources. In geographic analysis, the problem is often treated as one of “wombling” (see jacquez2003local; jacquez2003geographic; lu2005bayesian; li2011mining; ma2007bayesian; fitzpatrick2010, for some algorithmic approaches with applications). Probabilistic model-based approaches aiming to resolve this problem include, but are not limited to, developments in lu2007bayesian; ma2010hierarchical; li2015bayesian; banerjee2012bayesian; hanson2015spatial; leeBoundaryDetectionDisease2012; corpas-burgos2020serra and gaoSpatialDifferenceBoundary2023.
In pursuing a probabilistic learning mechanism for adjacency boundaries on a map from data available as aggregates over regions, as is typical of health administration data, we recognize that administrative boundaries between adjacent regions need not represent health outcomes measured there. In somewhat contrarian manner, the more conspicuous disease mapping models assume that neighboring regions are similar to each other. This runs counter to detecting health disparities, where we seek out differences and not similarities. Therefore, modeling spatial disparities needs to balance underlying spatial similarities based upon geographic proximity with the ability to detect differences among neighbors. li2015bayesian, gaoSpatialDifferenceBoundary2023 and aiello2023detecting, for example, propose classes of models that endow the spatial random effects with a discrete probability law using Dirichlet processes. This allows for nonzero probabilities on spatial random effects being equal to each, which are computed as posterior probabilities and used to classify spatial difference boundaries by controlling a Bayesian false discovery rate (on the lines of mullerFDRBayesianMultiple2006).
We aim to obviate some challenges seen in the aforementioned approaches. For one, the computation of the Bayesian nonparametric models devised for difference boundary analysis is cumbersome with the presence of weakly identifiable parameters that impede convergence of iterative estimation algorithms. A second issue is that, while semi-parametric approaches offer some flexibility, the hierarchical Dirichlet process obscures the spatial dependence between the random effects and makes the inference sensitive to values of hyper-parameters. The analysis presented in gaoSpatialDifferenceBoundary2023 also reveal that such models may not deliver optimal sensitivity and specificity with regard to detection of boundaries. Furthermore, ascertaining difference boundaries using such models take into account only the probability of spatial effects being different from each other and not the differences between such effects. How small differences among neighboring effects translate to higher values of the posterior probabilities of the differences remain unaccounted. And the methods do not offer a defense against the argument that small differences in the values of the spatial effects of adjacent regions need not constitute a disparity.
We exploit conjugate distribution theory for hierarchical models to account for the differences between the values of the spatial effects as well as the posterior probabilities of such differences exceeding a certain threshold. In this regard, we extend the approaches in lu2005bayesian and fitzpatrick2010, who use only posterior means or spatial residuals to detect differences, to full Bayesian inference using posterior distributions to quantify uncertainty. More specifically, we treat spatial disparities as a multiple comparison problem within a Bayesian inferential paradigm by testing for differences in spatial random effects between neighboring regions. Subsequently, we use posterior probabilities of such differences to detect spatial disparities while accounting for false discoveries with the Bayesian FDR (mullerOptimalSampleSize2004).
We provide an outline for the rest of the article. Section 2 discusses the problem of detecting differences between fixed effects in a linear model before we extend to spatial random effects. Section 3 develops the notion of a difference boundary and demonstrates a meaningful connection to traditional hypothesis testing in the simple linear regression model. Section LABEL:sec:simulation_experiments evaluates our methods using simulations over a map of U.S. counties followed by an analysis of tracheal, bronchus and lung cancer mortality data in Section LABEL:sec:application. Section LABEL:sec:_discussion concludes with a brief discussion.
2. Testing for Significant Contrasts in Bayesian Linear Regression
2.1. Conjugate Normal-Gamma Model
Consider the Bayesian linear regression model
$y=X\beta+\eta,\;\eta\sim N_{n}(0,\sigma^{2}V_{y})\;,$ (1)
where $y$ is an $n\times 1$ vector of observations, $X$ is an $n\times p$ matrix of predictors, $\beta$ is the unknown $p\times 1$ vector of corresponding regression coefficients, and $V_{y}$ is a $n\times n$ covariance matrix that is assumed to be completely specified by the analyst. A customary conjugate prior is
$\pi(\beta,\sigma^{2})=N(\beta\,|\,M_{0}m_{0},\sigma^{2}M_{0})\times\text{IG}(\sigma^{2}\,|\,a_{0},b_{0})\;,$ (2)
where $M_{0}$ is a $p\times p$ prior correlation matrix for $\beta$ that is fixed, $m_{0}$ is a fixed $p\times 1$ vector so that $\mathbb{E}[\beta]=M_{0}m_{0}$, $a_{0}$ and $b_{0}$ are known shape and rate parameters of the inverse-Gamma prior on $\sigma^{2}$. Familiar Bayesian calculations yield the joint posterior distribution
$\pi(\beta,\sigma^{2}\,|\,y)=N(\beta\,|\,Mm,\sigma^{2}M)\times\text{IG}(\sigma^{2}\,|\,a,b)\;,$ (3)
where $a=a_{0}+\frac{n}{2}$, $b=b_{0}+\frac{1}{2}\left(y^{\mathrm{\scriptscriptstyle T}}V_{y}^{-1}y+m_{0}^{\mathrm{\scriptscriptstyle T}}M_{0}m_{0}-m^{\mathrm{\scriptscriptstyle T}}Mm\right)$, $M=\left(M_{0}^{-1}+X^{\mathrm{\scriptscriptstyle T}}V_{y}^{-1}X\right)^{-1}$ and $m=m_{0}+X^{\mathrm{\scriptscriptstyle T}}V_{y}^{-1}y$. We study the problem of multiple pairwise comparison within the coefficients $\beta$ when $p>2$ and thus devote our attention to studying pairwise contrasts of the form $\beta_{i}-\beta_{j}$. The fundamental question is: after observing $y$, how can we make decisions about $\beta_{i}$ being different from $\beta_{j}$ for a large number of pairs $(i,j)$?
We must clarify how we intend to ascertain if $\beta_{i}$ is different from $\beta_{j}$. We are unable to simply use $\mathbb{P}(\beta_{i}=\beta_{j}\,|\,y)$ since $\mathbb{P}(\beta_{i}=\beta_{j}\,|\,y)=0$ for continuous models, such as (3), regardless of the observed data. Hence, it is not possible to directly evaluate the probability of $\beta_{i}=\beta_{j}$. One possibility is to use priors with point masses on the regression coefficients, such as slab-spike priors (mitchellBayesianVariableSelection1988) and several adaptations using mixture models to meaningfully evaluate $\mathbb{P}(\beta_{i}=0\,|\,y)$. For example, scottExplorationAspectsBayesian2006 investigate multiple comparisons of normal means to 00 by modeling each null hypothesis as independent Bernoulli trials with an unknown common failure probability. However, there is some difficulty in adopting a similar strategy for testing $\beta_{i}=\beta_{j}$ in the regression setting because posterior dependence among the elements of $\beta$ implies that it would be inappropriate to model the contrasts $\beta_{i}-\beta_{j}$ being zero as independent Bernoulli random variables. In other words, hypothesis testing of several contrasts on the regression coefficients is a multiple dependent hypothesis testing problem that prevents us from justifying a conditionally independent hierarchical model.
As it stands, the continuous conjugate prior under the standard Bayesian linear regression model is incompatible with difference probabilities that only consider exact equality. Changing the prior or the model sacrifices the analytical simplicity of the conjugate prior in the Bayesian linear regression setting, so we devote our attention to the third approach by considering a new class of difference probabilities. The goal here is to generate easily interpretable and informative statements about the parameters while also prioritizing model simplicity and computational cost.
2.2. Random Effects under an Augmented Linear System
The model in (1) can be criticized as overly simplistic if, for example, both spatial autocorrelation and unstructured random error substantially contribute to the true data generating mechanism. One model that attempts to address this is the BYM (Besag, York and Mollié) model, first introduced in besagBayesianImageRestoration1991, which includes both a spatially correlated and unstructured error component as part of a Poisson generalized linear model. However, simpsonPenalisingModelComponent2017 and rieblerIntuitiveBayesianSpatial2016 explain the error components are not identifiable and scaling issues complicate interpretation of the variance parameters in the BYM model. We follow the parametric form of the BYM model in rieblerIntuitiveBayesianSpatial2016 that ensures identifiable error components and imparts better interpretation. We write this BYM2 model as
$y=X\beta+\gamma+\eta,\;\;\gamma=\sigma\sqrt{\rho}\phi,\;\;\phi\sim N_{n}(0,V_{\phi}),\;\;\eta=\sigma\sqrt{1-\rho}v,\;\;v\sim N_{n}(0,I_{n})$ (4)
where $\phi$ is a $n\times 1$ vector of random effects, $\rho\in(0,1)$ controls the proportion of the total error attributed to the random effects’ variance structure, and $V_{\phi}$ is a known positive definite matrix. For subsequent analysis, we specify a conditionally autoregressive (CAR) structure, $V^{-1}_{\phi}=c(D_{W}-\alpha W)$, where $W$ is the adjacency matrix, $D_{W}$ is a diagonal matrix with the diagonal elements $D_{W,ii}$ being the number of neighbors of region $i$, $\alpha$ is the spatial smoothing parameter, and $c$ is a scaling factor. The parameter $\sigma^{2}$ captures the overall error variance if $c$ is set so that the geometric mean of the marginal prior variances is approximately one, and $\rho$ represents the proportion of total variance attributed to the spatial structure (rieblerIntuitiveBayesianSpatial2016).
If $\pi(\gamma,\beta,\sigma^{2},\rho)=N(\gamma\,|\,0,\sigma^{2}\rho V_{\phi})\times N(\beta\,|\,M_{0}m_{0},\sigma^{2}M_{0})\times\text{IG}(\sigma^{2}\,|\,a_{\sigma^{2}},b_{\sigma^{2}})\times\pi(\rho)$ is the prior for the parameters in (4), then it is helpful to express the joint posterior distribution in terms of the augmented model (banerjeeModelingMassiveSpatial2020a),
$\underbrace{\begin{bmatrix}y\\ M_{0}m_{0}\\ 0_{n}\end{bmatrix}}_{y_{\star}}=\underbrace{\begin{bmatrix}X&I_{n}\\ I_{p}&0_{p\times n}\\ 0_{n\times p}&I_{n}\end{bmatrix}}_{X_{\star}}\underbrace{\begin{bmatrix}\beta\\ \gamma\end{bmatrix}}_{\gamma_{\star}}+\underbrace{\begin{bmatrix}\eta\\ \eta_{\beta}\\ \eta_{\phi}\end{bmatrix}}_{\eta_{\star}},\;\;\eta_{\star}\sim N_{2n+p}\left(0_{2n+p},\sigma^{2}V_{y_{\star}}\right)\;,$ (5)
where $V_{y_{\star}}=\begin{bmatrix}(1-\rho)I_{n}&0&0\\ 0&M_{0}&0\\ 0&0&\rho V_{\phi}\end{bmatrix}$ and a prior $\pi(\gamma_{\star},\sigma^{2},\rho)=\text{IG}(\sigma^{2}\,|\,a_{\sigma^{2}},b_{\sigma^{2}})\times\pi(\rho)$ (since the prior for $\gamma_{\star}=(\beta^{\mathrm{\scriptscriptstyle T}},\gamma^{\mathrm{\scriptscriptstyle T}})^{\mathrm{\scriptscriptstyle T}}$ has been absorbed into the linear system). The resulting conditional posterior joint distribution, $\pi(\gamma_{\star},\sigma^{2}\,|\,y,\rho)$, is
${}\pi(\gamma_{\star},\sigma^{2}\,|\,y,\rho)=\text{N}_{n+p}\left(\gamma_{\star}\,|\,M_{\star}m_{\star},\sigma^{2}M_{\star}\right)\times\text{IG}\left(\sigma^{2}\,|\,a_{n},b_{n}\right)\;,$ (6)
where $M_{\star}=(X_{\star}^{\mathrm{\scriptscriptstyle T}}V_{y_{\star}}^{-1}X_{\star})^{-1}$, $m_{\star}=X_{\star}^{\mathrm{\scriptscriptstyle T}}V_{y_{\star}}^{-1}y_{\star}=\left(\frac{1}{1-\rho}y^{\mathrm{\scriptscriptstyle T}}X+m_{0}^{\mathrm{\scriptscriptstyle T}},\frac{1}{1-\rho}y^{\mathrm{\scriptscriptstyle T}}\right)^{\mathrm{\scriptscriptstyle T}}$, $a_{n}=a_{\sigma^{2}}+\frac{n}{2}$, and $b_{n}=b_{\sigma^{2}}+\frac{1}{2}\left(\frac{1}{1-\rho}y^{\mathrm{\scriptscriptstyle T}}y+m_{0}^{\mathrm{\scriptscriptstyle T}}M_{0}m_{0}-m_{\star}^{\mathrm{\scriptscriptstyle T}}M_{\star}m_{\star}\right)$. An alternate representation that is interpreted in terms of least squares regression expresses $M_{\star}m_{\star}=\widehat{\gamma}_{\star}$ as the general least square regression estimate solving $X_{\star}^{\mathrm{\scriptscriptstyle T}}V_{y_{\star}}^{-1}X_{\star}\widehat{\gamma}_{\star}=X_{\star}^{\mathrm{\scriptscriptstyle T}}V_{y_{\star}}^{-1}y_{\star}$ and the scale parameter $b_{n}=b_{\sigma^{2}}+\frac{ns^{2}}{2}$, where $s^{2}=\frac{1}{n}(y_{\star}-X_{\star}\widehat{\gamma}_{\star})^{\mathrm{\scriptscriptstyle T}}V_{y_{\star}}^{-1}(y_{\star}-X_{\star}\widehat{\gamma}_{\star})$ is the mean of the residual sums of squares for the linear model in (5).
If $\rho\in(0,1)$ is fixed, then we can draw exact samples from the joint posterior. Denote $A=X^{\mathrm{\scriptscriptstyle T}}X+(1-\rho)M_{0}^{-1}$ and $B=\frac{1-\rho}{\rho}V_{\phi}^{-1}+I_{n}-XA^{-1}X^{\mathrm{\scriptscriptstyle T}}$. Familiar calculations yield $\operatorname{Var}(\gamma\,|\,y,\sigma^{2},\rho)=\sigma^{2}(1-\rho)B^{-1}$, $\operatorname{Var}(\beta\,|\,y,\sigma^{2},\rho)=\sigma^{2}(1-\rho)\left(A^{-1}+A^{-1}X^{\mathrm{\scriptscriptstyle T}}B^{-1}XA^{-1}\right)$, and $\operatorname{Cov}(\beta,\gamma\,|\,y,\sigma^{2},\rho)=-\sigma^{2}(1-\rho)A^{-1}X^{\mathrm{\scriptscriptstyle T}}B^{-1}$. Similarly, from $\mathbb{E}[\gamma_{\star}\,|\,y,\sigma^{2},\rho]=M_{\star}m_{\star}$ we obtain $\mathbb{E}[\beta\,|\,y,\sigma^{2},\rho]=(1-\rho)/\rho\cdot A^{-1}X^{\mathrm{\scriptscriptstyle T}}B^{-1}V_{\phi}^{-1}y+(1-\rho)A^{-1}\left(X^{\mathrm{\scriptscriptstyle T}}B^{-1}X+I_{n}\right)m_{0}$ and $\mathbb{E}[\gamma\,|\,y,\sigma^{2},\rho]=B^{-1}(I_{n}-XA^{-1}X^{\mathrm{\scriptscriptstyle T}})y-(1-\rho)B^{-1}XA^{-1}m_{0}$. Since $\gamma=\sigma\sqrt{\rho}\phi$, we have
$\operatorname{Var}(\phi\,|\,y,\sigma^{2},\rho)=\left(V_{\phi}^{-1}+\frac{\rho}{1-\rho}\left(I_{n}-XA^{-1}X^{\mathrm{\scriptscriptstyle T}}\right)\right)^{-1}.$ (7)
A flat prior on $\beta$ considerably simplifies $\mathbb{E}[\gamma^{*}\,|\,y,\sigma^{2},\rho]$ and $\operatorname{Var}[\gamma^{*}\,|\,y,\sigma^{2},\rho]$. If $\pi(\gamma,\beta,\sigma^{2})=\text{N}_{n}(\gamma\,|\,0,\sigma^{2}\rho V_{\phi})\times\text{IG}(\sigma^{2}\,|\,a_{\sigma^{2}},b_{\sigma^{2}})$, then $B=\frac{1-\rho}{\rho}V_{\phi}^{-1}+I_{n}-H$l, where $H=X(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}$, $\mathbb{E}[\beta\,|\,y,\sigma^{2},\rho]=(1-\rho)\cdot(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}B^{-1}V_{\phi}^{-1}y/\rho$, $\operatorname{Var}(\beta\,|\,y,\sigma^{2},\rho)=\sigma^{2}(1-\rho)(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}+\sigma^{2}(1-\rho)(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}B^{-1}X(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}$, $\mathbb{E}[\gamma\,|\,y,\sigma^{2},\rho]=B^{-1}(I-H)y$, $\operatorname{Var}(\gamma\,|\,y,\sigma^{2},\rho)=\sigma^{2}(1-\rho)B^{-1}$, and $\operatorname{Cov}(\beta,\gamma\,|\,y,\sigma^{2},\rho)=-\sigma^{2}(1-\rho)(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}B^{-1}$. Since $\gamma=\sigma\sqrt{\rho}\phi$, we obtain
$\operatorname{Var}(\phi\,|\,y,\sigma^{2},\rho)=\left(V_{\phi}^{-1}+\frac{\rho}{1-\rho}\left(I_{n}-X(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}\right)\right)^{-1}\;.$ (8)
In (4), inference on differences within the fixed effects $\beta$ and the spatial residuals $\gamma$ are equivalent problems if $\rho$ is known. This implies that a decision framework for the Bayesian fixed effects model specified in (1) can be extended to draw inference on both fixed and random effects in the model given by (4) after conditioning on $\rho$ and $V_{\phi}$. Thus, we turn to the original problem of computing difference probabilities in the fixed effects model before extending to spatial disparities.
2.3. Limiting cases
The limits of the conditional posterior distribution of $(\beta,\gamma)$ as $\rho\rightarrow 0^{+}$ or as $\rho\rightarrow 1^{-}$ can be analytically derived when the prior is non-informative on $\beta$. By the spectral decomposition, $V^{1/2}_{\phi}(I-H)V^{1/2}_{\phi}=CDC^{\mathrm{\scriptscriptstyle T}}$ where $C$ is an orthogonal matrix and $D=\text{diag}(d_{1},\ldots,d_{n})$ with entries that are the non-negative eigenvalues of $V_{\phi}^{1/2}\left(I_{n}-H\right)V_{\phi}^{1/2}$. Let $U=V^{1/2}_{\phi}C$, then $U$ is invertible, $U^{\mathrm{\scriptscriptstyle T}}V^{-1}_{\phi}U=I_{n}$, $U^{\mathrm{\scriptscriptstyle T}}(I_{n}-H)U=D$, $(1-\rho)B^{-1}=U\left(\frac{1}{\rho}I_{n}+\frac{1}{1-\rho}D\right)^{-1}U^{\mathrm{\scriptscriptstyle T}}$, and $B^{-1}(I-H)=U\left(\frac{1-\rho}{\rho}I_{n}+D\right)^{-1}DU^{-1}$. Here, $U$ does not depend on $\rho$ and $\left(\frac{1}{\rho}I_{n}+\frac{1}{1-\rho}D\right)^{-1}$ is diagonal with elements $\left(\frac{1}{\rho}I_{n}+\frac{1}{1-\rho}D\right)^{-1}_{ii}=\left(\frac{1}{\rho}+\frac{d_{i}}{1-\rho}\right)^{-1}$and we immediately obtain $\lim_{\rho\rightarrow 0^{+}}\left(\frac{1}{\rho}I_{n}+\frac{1}{1-\rho}D\right)^{-1}=0_{n\times n}$. Next, $B^{-1}(I-H)=U\left(\frac{1-\rho}{\rho}I_{n}+D\right)^{-1}DU^{-1}$, where $\left(\left(\frac{1-\rho}{\rho}I_{n}+D\right)^{-1}D\right)_{ii}=d_{i}\left(\frac{1-\rho}{\rho}+d_{i}\right)^{-1}$ which is 0 for any $\rho\in(0,1)$ if $d_{i}=0$, and converges to 0 as $\rho\rightarrow 0^{+}$ if $d_{i}>0$. Thus, $\lim_{\rho\rightarrow 0^{+}}\left(\left(\frac{1-\rho}{\rho}I_{n}+D\right)^{-1}D\right)=0_{n\times n}$. These imply that if the prior is non-informative on $\beta$, then $\lim_{\rho\rightarrow 0^{+}}\pi(\beta,\gamma,\sigma^{2}\,|\,y,\rho)$ collapses to the degenerate mass at $\gamma=0_{n}$.
Turning to $\rho\rightarrow 1^{-}$, we note that if $d_{i}=0$, $\left(\frac{1}{\rho}I_{n}+\frac{1}{1-\rho}D\right)^{-1}_{ii}=\rho$, which converges to $1$ as $\rho\rightarrow 1^{-}$. If $d_{i}>0$, then $\lim_{\rho\rightarrow 1^{-}}\left(\frac{1}{\rho}+\frac{d_{i}}{1-\rho}\right)^{-1}=0$ and $\lim_{\rho\rightarrow 1^{-}}d_{i}\left(\frac{1-\rho}{\rho}+d_{i}\right)^{-1}=1$. Thus, $\lim_{\rho\rightarrow 1^{-}}\left(\frac{1}{\rho}I_{n}+\frac{1}{1-\rho}D\right)^{-1}=I_{n}-D^{\star}$ and $\lim_{\rho\rightarrow 1^{-}}\left(\frac{1-\rho}{\rho}I_{n}+D\right)^{-1}D=D^{\star}$ where $D^{\star}\in\mathbb{R}^{n\times n}$ is diagonal with $D_{ii}^{\star}=I(d_{i}>0)$ for $i=1,\ldots,n$. Further simplifications reveal that $\lim_{\rho\rightarrow 1^{-}}\mathbb{E}[\beta\,|\,y,\sigma^{2},\rho]=(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}\left(y-UD^{\star}U^{-1}e\right)$, and $\lim_{\rho\rightarrow 1^{-}}\mathbb{E}[\gamma\,|\,y,\sigma^{2},\rho]=UD^{\star}U^{-1}e$, where $e=(I-H)y$. Second order moments have limits given by $\lim_{\rho\rightarrow 1^{-}}\operatorname{Var}(\beta\,|\,y,\sigma^{2},\rho)=\sigma^{2}(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}VX(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}$, $\lim_{\rho\rightarrow 1^{-}}\operatorname{Var}(\gamma\,|\,y,\sigma^{2},\rho)=\sigma^{2}V$, and $\lim_{\rho\rightarrow 1^{-}}\operatorname{Cov}(\beta,\gamma\,|\,y,\sigma^{2},\rho)=-\sigma^{2}(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}V$ where $V=V_{\phi}-UD^{\star}U^{\mathrm{\scriptscriptstyle T}}$. Furthermore, $\lim_{\rho\rightarrow 1^{-}}b_{n}=b_{\sigma^{2}}+\frac{1}{2}v^{\mathrm{\scriptscriptstyle T}}D^{\star}v$, where $v=U^{-1}e$ (see Section LABEL:sec:DIC_analysis). Therefore, $\lim_{\rho\rightarrow 1^{-}}\pi(\sigma^{2}\,|\,y,\rho)=\text{IG}(\sigma^{2}\,|\,a_{\sigma^{2}}+n/2,b_{\sigma^{2}}+v^{\mathrm{\scriptscriptstyle T}}D^{\star}v/2)$.
3. A New Framework for Detecting Spatial Disparities
3.1. Connection to Classical Hypothesis Testing
Whereas $\mathbb{P}(c_{k}^{\mathrm{\scriptscriptstyle T}}\beta=0\,|\,y)=0$ for a continuous posterior distribution, we can test a point null hypotheses $H^{(k)}_{0}:c_{k}^{\mathrm{\scriptscriptstyle T}}\beta=0$ against $H^{(k)}_{1}:c_{k}^{\mathrm{\scriptscriptstyle T}}\beta\neq 0$ for $K$ non-zero $p\times 1$ vectors $c_{1},\ldots,c_{K}\in\mathbb{R}^{p}$ using the classical linear model, $Y=X\beta+\eta,\eta\sim N_{n}(0_{n},\sigma^{2}V_{y})$, where $\beta$ is a fixed but unknown $p\times 1$ vector (see, e.g., seberHypothesisTesting2003). For each $c_{k}$, this goal is accomplished by computing the test statistic $t_{\text{obs}}^{(k)}=\frac{c_{k}^{\mathrm{\scriptscriptstyle T}}\widehat{\beta}}{\widehat{\sigma}\sqrt{c_{k}^{\mathrm{\scriptscriptstyle T}}(X^{\mathrm{\scriptscriptstyle T}}V_{y}^{-1}X)^{-1}c_{k}}}$, where $\widehat{\beta}=(X^{\mathrm{\scriptscriptstyle T}}V_{y}^{-1}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}V_{y}^{-1}y$ and $\widehat{\sigma}^{2}=\frac{1}{n-p}\left(y-X\widehat{\beta}\right)^{\mathrm{\scriptscriptstyle T}}V_{y}^{-1}\left(y-X\widehat{\beta}\right)$ are the generalized least squares estimates. The classical approach computes a p-value $p_{k}=\mathbb{P}\left(T>\left|t^{(k)}_{\text{obs}}\right|\right)$ where $T\sim t_{n-p}$ and accounts for multiple testing by computing adjusted p-values. If the adjusted p-value is below a pre-specified significance threshold, then $H^{(k)}_{0}$ is rejected.
In the Bayesian framework, which we adopt here, we maintain the continuous conjugate prior specified by (2) and recast the decision problem to consider the posterior probabilities of absolute standardized linear combinations among elements of $\beta$ being greater than a threshold $\epsilon$. For all $k=1,\ldots,K$ and any $\epsilon>0$, we define
$v_{k}(\epsilon)=\mathbb{P}\left(\left.\frac{|c_{k}^{\mathrm{\scriptscriptstyle T}}\beta|}{\sigma\sqrt{c_{k}^{\mathrm{\scriptscriptstyle T}}Mc_{k}}}>\epsilon\;\right|\;y\right).$ (9)
We refer to these posterior probabilities as $\epsilon$-difference probabilities, where $\epsilon$ represents a universal threshold across all $K$ linear combinations and represents the number of posterior standard deviation units that $|c_{k}^{\mathrm{\scriptscriptstyle T}}\beta|$ is away from zero. If $\beta^{(1)},\ldots,\beta^{(N)}$ and $\sigma^{(1)},\ldots,\sigma^{(N)}$ are $N$ samples drawn from their respective posterior distributions, then we estimate (9) using
$\widehat{v}_{k}(\epsilon)=\frac{1}{N}\sum_{t=1}^{N}\text{I}\left(\frac{|c_{k}^{\mathrm{\scriptscriptstyle T}}\beta^{(t)}|}{\sigma^{(t)}\sqrt{c_{k}^{\mathrm{\scriptscriptstyle T}}Mc_{k}}}>\epsilon\right)$ (10)
for any given $\epsilon$. Using a conjugate prior, we efficiently obtain exact posterior samples of $\beta\,|\,y$ using (3). Specific domain applications may provide further insights into possible constraints on $\epsilon$ and what is considered to be a clinically meaningful difference between coefficients $\beta_{i}$ and $\beta_{j}$, which may, for example, represent mean outcomes for different treatment groups in an ANOVA.
Classical p-values and the probabilities $\left\{v_{k}(\epsilon)\right\}_{k}$ defined in (9) have opposite interpretation: a small p-value, say $p_{k}$, suggests $|c_{k}^{\mathrm{\scriptscriptstyle T}}\beta|$ is far from 0, while small $v_{k}(\epsilon)$ suggests $|c_{k}^{\mathrm{\scriptscriptstyle T}}\beta|$ is likely to be close to 0. The posterior distribution of $\beta$ with a flat prior is connected to the generalized least squares estimate $\widehat{\beta}$ in the Bayesian linear regression model. Hence, with a specific prior, the p-values and $\epsilon$-difference probabilities are perfectly correlated.
Theorem 1.
Let $c_{1},\ldots,c_{K}\in\mathbb{R}^{p}$, $X\in\mathbb{R}^{n\times p}$, $y\in\mathbb{R}^{n}$, and $V_{y}\in\mathbb{R}^{n\times n}$ be fixed. Under the standard linear model $y\sim N_{n}(X\beta,\sigma^{2}V_{y})$, let $p_{k}$ be the p-value for testing the null hypothesis $H_{0}^{(k)}:c_{k}^{\mathrm{\scriptscriptstyle T}}\beta=0$ against $H_{1}^{(k)}:c_{k}^{\mathrm{\scriptscriptstyle T}}\beta\neq 0$ for $k=1,\ldots,K$. Then, $p_{k}>p_{k^{\prime}}$ if and only if $v_{k}(\epsilon)<v_{k^{\prime}}(\epsilon)$ for all $k\neq k^{\prime}$, where $v_{k}(\epsilon)=\mathbb{P}\left(\left.\frac{|c_{k}^{\mathrm{\scriptscriptstyle T}}\beta|}{\sigma\sqrt{c_{k}^{\mathrm{\scriptscriptstyle T}}Mc_{k}}}>\epsilon\;\right|\;y\right)$ under $\pi(\beta,\sigma^{2})\propto\pi(\sigma^{2})$ for any given $\epsilon>0$.
Proof.
See Appendix LABEL:append:prop1_proof.
For the linear model, the Bayesian approach of ranking $\epsilon$-difference probabilities can replicate the rankings of the classical p-values by using a flat $\beta$ prior. Theorem 1 implies that the ascending order of the p-values associated with null hypotheses of the form $H_{0}^{(k)}:c_{k}^{\mathrm{\scriptscriptstyle T}}\beta=0$ with two-sided alternative $H_{1}^{(k)}:c_{k}^{\mathrm{\scriptscriptstyle T}}\beta\neq 0$ is equivalent to the descending order of the Bayesian $\epsilon$-difference probabilities. Since this result holds for all $\epsilon>0$ and the classical procedure does not depend on $\epsilon$, the Bayesian rankings are stable with respect to $\epsilon$.
3.2. Extension to Spatial Random Effects
We next consider the model in (4), where each observation corresponds to one areal region. We treat $V_{\phi}$ as a fixed random effects variance structure, whereas $\rho$, which controls the proportion of the total error variance attributed to the structured component, can either be a fixed hyperparameter or a random variable with a prior density. Generalizing the idea of $\epsilon$-difference probabilities in the fixed effects model, we now conduct posterior inference on random variables of the form $r_{k}(\epsilon)=\text{I}\left(\frac{|c_{k}^{\mathrm{\scriptscriptstyle T}}\gamma_{\star}|}{\sigma\sqrt{c_{k}^{\mathrm{\scriptscriptstyle T}}M_{\star}c_{k}}}>\epsilon\right)$, for any given $\epsilon>0$. For $k=1\ldots,K$, we define the conditional posterior probability
$h_{k}(\epsilon;\rho)=\mathbb{P}\left(\left.\frac{|c_{k}^{\mathrm{\scriptscriptstyle T}}\gamma_{\star}|}{\sigma\sqrt{c_{k}^{\mathrm{\scriptscriptstyle T}}M_{\star}c_{k}}}>\epsilon\;\right|\;y,\rho\right),$ (11)
where $c_{k}\in\mathbb{R}^{n+p}$ is typically a contrast on $\gamma_{\star}=(\beta^{\mathrm{\scriptscriptstyle T}},\gamma^{\mathrm{\scriptscriptstyle T}})^{\mathrm{\scriptscriptstyle T}}$. If $c_{k}^{\mathrm{\scriptscriptstyle T}}\gamma_{\star}=a_{k}^{\mathrm{\scriptscriptstyle T}}\gamma$ for some $a_{k}\in\mathbb{R}^{n}$, then $h_{k}(\epsilon;\rho)=\mathbb{P}\left(\left.\frac{|a_{k}^{\mathrm{\scriptscriptstyle T}}\phi|}{\sqrt{a_{k}^{\mathrm{\scriptscriptstyle T}}\operatorname{Var}(\phi\,|\,y,\sigma^{2},\rho)a_{k}}}>\epsilon\;\right|\;y,\rho\right)$ where $\operatorname{Var}(\phi\,|\,y,\sigma^{2},\rho)$ is given by (7).
In the context of detecting spatial disparities, consider a set of $K$ pairs of neighboring regions denoted as $L=\{(i,j):i<j,i\sim j\}$ where $i\sim j$ means that regions $i$ and $j$ are neighbors. If $c_{ij}\in\mathbb{R}^{n+p}$ and $a_{ij}\in\mathbb{R}^{n}$ such that $c_{ij}^{\mathrm{\scriptscriptstyle T}}\gamma_{\star}=a_{ij}^{\mathrm{\scriptscriptstyle T}}\gamma=\gamma_{i}-\gamma_{j}$, then we denote the corresponding $\epsilon$-difference probability as $h_{ij}(\epsilon;\rho)=\mathbb{P}\left(\left.\frac{|\phi_{i}-\phi_{j}|}{\sqrt{a_{ij}^{\mathrm{\scriptscriptstyle T}}\operatorname{Var}(\phi\,|\,y,\sigma^{2},\rho)a_{ij}}}>\epsilon\;\right|\;y,\rho\right)$. The set $\left\{h_{ij}(\epsilon;\rho)\right\}_{(i,j)\in L}$ then represents the $\epsilon$-difference probabilities of all boundaries under consideration.
According to the equivalence between the model given by (4) and the augmented linear system discussed in Section 2.2, similar insight to the proof of Theorem 1 reveals that the rankings of the conditional posterior probabilities $h_{1}(\epsilon),\ldots,h_{K}(\epsilon)$ are stable with respect to $\epsilon$ for any $\rho\in(0,1)$.