### v3

# supplementary

Abstract.
Epidemiological investigations of regionally aggregated spatial data often involve detecting spatial health disparities among neighboring regions on a map of disease mortality or incidence rates. Analyzing such data introduces spatial dependence among the health outcomes and seeks to report statistically significant spatial disparities by delineating boundaries that separate neighboring regions with disparate health outcomes. However, there are statistical challenges to appropriately defining what constitutes a spatial disparity and to construct robust probabilistic inference for spatial disparities. We enrich the familiar Bayesian linear regression framework to introduce spatial autoregression and offer model-based detection of spatial disparities. We derive exploitable analytical tractability that considerably accelerates computation. Simulation experiments conducted over a county map of the entire United States demonstrate the effectiveness of our method and we apply our method to a data set from the Institute of Health Metrics and Evaluation (IHME) on age-standardized US county-level estimates of lung cancer mortality rates.
1. Introduction
We develop statistical methods for detecting spatial disparities on maps. Disparities in health outcomes is a rapidly evolving field within public health (see, e.g., rao2023) and spatial variation in health outcomes affects how disparities manifest on disease maps. Spatial data analysis in epidemiological investigations is extensively documented (see, e.g., waller2004applied; waller2010handbook; lawson2013statistical; lawson2016handbook, and references therein). Inference on disparities proceeds from “difference boundaries” that represent significant differences among neighbors and delineate regions with significantly different disease mortality and incidence rates, which assists in improved geographic allocation of health care resources. In geographic analysis, the problem is often called “wombling” (see jacquez2003local; jacquez2003geographic; lu2005bayesian; li2011mining; ma2007bayesian; fitzpatrick2010, for some algorithmic approaches with applications). Probabilistic model-based approaches include developments in lu2007bayesian; ma2010hierarchical; li2015bayesian; banerjee2012bayesian; hanson2015spatial; leeBoundaryDetectionDisease2012; corpas-burgos2020serra; gaoSpatialDifferenceBoundary2023 and aiello2023detecting.
While the aforementioned articles present an impressive range of methods, they are broadly classified into two groups. The first embraces hierarchical models and Bayesian inference for the spatial adjacency matrix by modeling its elements. The second endows the spatial random effects with a probability law that accommodates nonzero probabilities on neighboring spatial random effects being equal (often using an areal Dirichlet process or some adaptation thereof), which are computed as posterior probabilities. Both groups introduce spatial dependence and aim to balance underlying spatial similarities with the ability to detect differences among neighbors. Our current contribution offers a computationally scalable and analytically tractable approach to this problem. Instead of modeling spatial effects as realizations of a Dirichlet process, which struggles to learn from the data and impedes convergence of iterative estimation algorithms, we devise fully model-based Bayesian inference within an augmented spatial linear regression framework. More specifically, we treat spatial disparities as a multiple comparison problem within a Bayesian inferential paradigm by testing for differences in spatial random effects between neighboring regions. Subsequently, we use posterior probabilities of such differences to detect spatial disparities while controlling a Bayesian false discovery rate or FDR (mullerFDRBayesianMultiple2006). Benefits of our approach include clearer interpretation of spatial effects, computational simplicity, accounting for small differences in the values of neighboring spatial effects of adjacent regions in ascertaining spatial disparity, and delivering improved sensitivity and specificity for difference boundary detection (as seen in the analysis by gaoSpatialDifferenceBoundary2023).
We exploit closed form distribution theory to account for the differences between the values of the spatial effects and evaluate posterior probabilities of such differences exceeding a certain threshold. We extend lu2005bayesian and fitzpatrick2010, who use only posterior means or spatial residuals to detect differences, to fully model-based inference. We provide an outline for the rest of the article. Section 2 discusses the problem of detecting differences between fixed effects in a linear model before we extend to spatial random effects. Section 3 develops the notion of a difference boundary, demonstrates a meaningful connection to traditional hypothesis testing in the simple linear regression model, and considers extensions to a generalized linear model. Section LABEL:sec:simulation_experiments evaluates our methods using simulated data over a map of U.S. counties followed by an analysis of lung cancer mortality data in Section LABEL:sec:application. Section LABEL:sec:_discussion concludes with a brief discussion.
2. Assessing Contrasts in Bayesian Linear Regression
2.1. Random Effects under an Augmented Linear System
We follow rieblerIntuitiveBayesianSpatial2016 who adapted the “BYM” model (besagBayesianImageRestoration1991) to ensure identifiable parameters and afford better interpretation. We write this “BYM2” model as
$y=X\beta+\gamma+\eta,\;\;\gamma=\sigma\sqrt{\rho}\phi,\;\;\phi\sim N_{n}(0,V_{\phi}),\;\;\eta=\sigma\sqrt{1-\rho}v,\;\;v\sim N_{n}(0,I_{n})$ (1)
where $\phi$ is an $n\times 1$ vector of random effects, $V_{\phi}$ is a known positive definite matrix. For subsequent analysis, we specify a conditionally autoregressive (CAR) structure, $V^{-1}_{\phi}=c(D_{W}-\alpha W)$, where $W$ is the adjacency matrix, $D_{W}$ is diagonal with the diagonal elements $D_{W,ii}$ being the number of neighbors of region $i$, $\alpha$ is the spatial smoothing parameter, and $c$ is a scaling factor. The parameter $\sigma^{2}$ captures the overall error variance if $c$ is set so that the geometric mean of the marginal prior variances is approximately one, and $\rho$ represents the proportion of total variance attributed to the spatial structure (rieblerIntuitiveBayesianSpatial2016).
If $\pi(\gamma,\beta,\sigma^{2},\rho)=N(\gamma\,|\,0,\sigma^{2}\rho V_{\phi})\times N(\beta\,|\,M_{0}m_{0},\sigma^{2}M_{0})\times\text{IG}(\sigma^{2}\,|\,a_{\sigma^{2}},b_{\sigma^{2}})\times\pi(\rho)$ is the prior for (1), then the joint distribution arises from the augmented linear model,
$\underbrace{\begin{bmatrix}y\\ M_{0}m_{0}\\ 0_{n}\end{bmatrix}}_{y_{\star}}=\underbrace{\begin{bmatrix}X&I_{n}\\ I_{p}&0_{p\times n}\\ 0_{n\times p}&I_{n}\end{bmatrix}}_{X_{\star}}\underbrace{\begin{bmatrix}\beta\\ \gamma\end{bmatrix}}_{\gamma_{\star}}+\underbrace{\begin{bmatrix}\eta\\ \eta_{\beta}\\ \eta_{\phi}\end{bmatrix}}_{\eta_{\star}},\;\;\eta_{\star}\sim N_{2n+p}\left(0_{2n+p},\sigma^{2}V_{y_{\star}}\right)\;,$ (2)
where $V_{y_{\star}}=\begin{bmatrix}(1-\rho)I_{n}&0&0\\ 0&M_{0}&0\\ 0&0&\rho V_{\phi}\end{bmatrix}$ and a prior $\pi(\gamma_{\star},\sigma^{2},\rho)=\text{IG}(\sigma^{2}\,|\,a_{\sigma^{2}},b_{\sigma^{2}})\times\pi(\rho)$ (since the prior for $\gamma_{\star}=(\beta^{\mathrm{\scriptscriptstyle T}},\gamma^{\mathrm{\scriptscriptstyle T}})^{\mathrm{\scriptscriptstyle T}}$ has been absorbed into the linear system). The resulting conditional posterior joint distribution, $\pi(\gamma_{\star},\sigma^{2}\,|\,y,\rho)$, is
${}\pi(\gamma_{\star},\sigma^{2}\,|\,y,\rho)=\text{N}_{n+p}\left(\gamma_{\star}\,|\,M_{\star}m_{\star},\sigma^{2}M_{\star}\right)\times\text{IG}\left(\sigma^{2}\,|\,a_{n},b_{n}\right)\;,$ (3)
where $M_{\star}=(X_{\star}^{\mathrm{\scriptscriptstyle T}}V_{y_{\star}}^{-1}X_{\star})^{-1}$, $m_{\star}=X_{\star}^{\mathrm{\scriptscriptstyle T}}V_{y_{\star}}^{-1}y_{\star}=\left(\frac{1}{1-\rho}y^{\mathrm{\scriptscriptstyle T}}X+m_{0}^{\mathrm{\scriptscriptstyle T}},\frac{1}{1-\rho}y^{\mathrm{\scriptscriptstyle T}}\right)^{\mathrm{\scriptscriptstyle T}}$, $a_{n}=a_{\sigma^{2}}+\frac{n}{2}$, and $b_{n}=b_{\sigma^{2}}+\frac{1}{2}\left(\frac{1}{1-\rho}y^{\mathrm{\scriptscriptstyle T}}y+m_{0}^{\mathrm{\scriptscriptstyle T}}M_{0}m_{0}-m_{\star}^{\mathrm{\scriptscriptstyle T}}M_{\star}m_{\star}\right)$. The posterior mean $M_{\star}m_{\star}=\widehat{\gamma}_{\star}$ is the general least square regression estimate solving $X_{\star}^{\mathrm{\scriptscriptstyle T}}V_{y_{\star}}^{-1}X_{\star}\widehat{\gamma}_{\star}=X_{\star}^{\mathrm{\scriptscriptstyle T}}V_{y_{\star}}^{-1}y_{\star}$ and $b_{n}=b_{\sigma^{2}}+\frac{ns^{2}}{2}$, where $s^{2}=\frac{1}{n}(y_{\star}-X_{\star}\widehat{\gamma}_{\star})^{\mathrm{\scriptscriptstyle T}}V_{y_{\star}}^{-1}(y_{\star}-X_{\star}\widehat{\gamma}_{\star})$ is the mean residual sums of squares from (2).
If $\rho\in(0,1)$ is fixed, then we can draw exact posterior samples from (3). Noting that $\gamma=\sigma\sqrt{\rho}\phi$ and carrying out some familiar linear algebra calculations, we obtain
$\operatorname{Var}(\phi\,|\,y,\sigma^{2},\rho)=\left(V_{\phi}^{-1}+\frac{\rho}{1-\rho}\left(I_{n}-XA^{-1}X^{\mathrm{\scriptscriptstyle T}}\right)\right)^{-1},$ (4)
where $A=X^{\mathrm{\scriptscriptstyle T}}X+(1-\rho)M_{0}^{-1}$. A flat prior on $\beta$, i.e., $M_{0}^{-1}=O$, yields $A=X^{\mathrm{\scriptscriptstyle T}}X$ and
$\lim_{M_{0}^{-1}\to O}\operatorname{Var}(\phi\,|\,y,\sigma^{2},\rho)=\left(V_{\phi}^{-1}+\frac{\rho}{1-\rho}\left(I_{n}-H\right)\right)^{-1},\;\mbox{ where }\;H=X(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}\;.$ (5)
These expressions are used in Section 3.2 to formulate difference boundary detection.
2.2. Limiting cases
The limits of the conditional posterior distribution of $(\beta,\gamma)$ as $\rho\rightarrow 0^{+}$ or as $\rho\rightarrow 1^{-}$ are analytically derived when the prior is non-informative on $\beta$. By the spectral decomposition, $V^{1/2}_{\phi}(I-H)V^{1/2}_{\phi}=CDC^{\mathrm{\scriptscriptstyle T}}$ where $C$ is an orthogonal matrix and $D=\text{diag}(d_{1},\ldots,d_{n})$ with entries that are the non-negative eigenvalues of $V_{\phi}^{1/2}\left(I_{n}-H\right)V_{\phi}^{1/2}$. Let $U=V^{1/2}_{\phi}C$, then $U$ is invertible, $U^{\mathrm{\scriptscriptstyle T}}V^{-1}_{\phi}U=I_{n}$, $U^{\mathrm{\scriptscriptstyle T}}(I_{n}-H)U=D$, $\operatorname{Var}(\gamma\,|\,y,\sigma^{2},\rho)=\sigma^{2}U\left(\frac{1}{\rho}I_{n}+\frac{1}{1-\rho}D\right)^{-1}U^{\mathrm{\scriptscriptstyle T}}$, and $\mathbb{E}[\gamma\,|\,y,\sigma^{2},\rho]=U\left(\frac{1-\rho}{\rho}I_{n}+D\right)^{-1}DU^{-1}y$. Calculations reveal $\lim_{\rho\rightarrow 0^{+}}\operatorname{Var}(\gamma\,|\,y,\sigma^{2},\rho)=0_{n\times n}$ and $\lim_{\rho\rightarrow 0^{+}}\mathbb{E}[\gamma\,|\,y,\sigma^{2},\rho]=0_{n}$. Therefore, if the prior on $\beta$ is flat, then $\lim_{\rho\rightarrow 0^{+}}\pi(\beta,\gamma,\sigma^{2}\,|\,y,\rho)$ collapses to the degenerate mass at $\gamma=0_{n}$.
We also obtain $\lim_{\rho\rightarrow 1^{-}}\left(\frac{1}{\rho}I_{n}+\frac{1}{1-\rho}D\right)^{-1}=I_{n}-D^{\star}$ and $\lim_{\rho\rightarrow 1^{-}}\left(\frac{1-\rho}{\rho}I_{n}+D\right)^{-1}D=D^{\star}$ where $D^{\star}\in\mathbb{R}^{n\times n}$ is diagonal with $D_{ii}^{\star}=I(d_{i}>0)$ for $i=1,\ldots,n$. Further simplifications reveal that $\mathbb{E}[\beta\,|\,y,\sigma^{2},\rho]\rightarrow(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}\left(y-UD^{\star}U^{-1}e\right)$ and $\mathbb{E}[\gamma\,|\,y,\sigma^{2},\rho]\rightarrow UD^{\star}U^{-1}e$ as $\rho\rightarrow 1^{-}$, where $e=(I-H)y$. Furthermore, $\operatorname{Var}(\beta\,|\,y,\sigma^{2},\rho)\rightarrow\sigma^{2}(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}VX(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}$, $\operatorname{Var}(\gamma\,|\,y,\sigma^{2},\rho)\rightarrow\sigma^{2}V$ and $\operatorname{Cov}(\beta,\gamma\,|\,y,\sigma^{2},\rho)\rightarrow-\sigma^{2}(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}V$ as $\rho\rightarrow 1^{-}$, where $V=V_{\phi}-UD^{\star}U^{\mathrm{\scriptscriptstyle T}}$. Finally, $\lim_{\rho\rightarrow 1^{-}}b_{n}=b_{\sigma^{2}}+\frac{1}{2}v^{\mathrm{\scriptscriptstyle T}}D^{\star}v$, where $v=U^{-1}e$ (see Section LABEL:sec:DIC_analysis). Therefore, $\lim_{\rho\rightarrow 1^{-}}\pi(\sigma^{2}\,|\,y,\rho)=\text{IG}(\sigma^{2}\,|\,a_{\sigma^{2}}+n/2,b_{\sigma^{2}}+v^{\mathrm{\scriptscriptstyle T}}D^{\star}v/2)$. These limiting cases have implications for subsequent boundary detection (see Section LABEL:subsec:epsilon_intro).
3. A New Framework for Detecting Spatial Disparities
3.1. Connection to Classical Hypothesis Testing
We first briefly address testing for contrasts in Bayesian linear regression without spatial effects, i.e., $\gamma=0$ in (1). Whereas $\mathbb{P}(c_{k}^{\mathrm{\scriptscriptstyle T}}\beta=0\,|\,y)=0$ for a continuous posterior distribution, we can test $H^{(k)}_{0}:c_{k}^{\mathrm{\scriptscriptstyle T}}\beta=0$ against $H^{(k)}_{1}:c_{k}^{\mathrm{\scriptscriptstyle T}}\beta\neq 0$ for $K$ non-zero $p\times 1$ vectors $c_{1},\ldots,c_{K}$ using a classical linear model, $y=X\beta+\eta,\eta\sim N_{n}(0_{n},\sigma^{2}V_{y})$, where $\beta$ is fixed but unknown (e.g., seberHypothesisTesting2003). We compute $t_{\text{obs}}^{(k)}=\frac{c_{k}^{\mathrm{\scriptscriptstyle T}}\widehat{\beta}}{\widehat{\sigma}\sqrt{c_{k}^{\mathrm{\scriptscriptstyle T}}(X^{\mathrm{\scriptscriptstyle T}}V_{y}^{-1}X)^{-1}c_{k}}}$ for each $c_{k}$, where $\widehat{\beta}=(X^{\mathrm{\scriptscriptstyle T}}V_{y}^{-1}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}V_{y}^{-1}y$ and $(n-p)\widehat{\sigma}^{2}=\left(y-X\widehat{\beta}\right)^{\mathrm{\scriptscriptstyle T}}V_{y}^{-1}\left(y-X\widehat{\beta}\right)$ are the generalized least squares estimates. We reject $H^{(k)}_{0}$ if the p-values $p_{k}=\mathbb{P}\left(T>\left|t^{(k)}_{\text{obs}}\right|\right)$, where $T\sim t_{n-p}$, are below a multiplicity adjusted threshold.
For Bayesian inference, we retain the continuous conjugate prior on $\beta$ and evaluate posterior probabilities for the absolute standardized linear combinations among elements of $\beta$ to exceed a threshold $\epsilon$. Hence, for all $k=1,\ldots,K$ and any $\epsilon>0$, we define
$v_{k}(\epsilon)=\mathbb{P}\left(\left.\frac{|c_{k}^{\mathrm{\scriptscriptstyle T}}\beta|}{\sigma\sqrt{c_{k}^{\mathrm{\scriptscriptstyle T}}Mc_{k}}}>\epsilon\;\right|\;y\right).$ (6)
We refer to these posterior probabilities as $\epsilon$-difference probabilities, where $\epsilon$ represents a universal threshold across all $K$ linear combinations indicating the number of posterior standard deviation units that $|c_{k}^{\mathrm{\scriptscriptstyle T}}\beta|$ is away from 00. If $\beta^{(1)},\ldots,\beta^{(N)}$ and $\sigma^{(1)},\ldots,\sigma^{(N)}$ are $N$ samples drawn from their respective posterior distributions, then we estimate (6) using
$\widehat{v}_{k}(\epsilon)=\frac{1}{N}\sum_{t=1}^{N}\text{I}\left(\frac{|c_{k}^{\mathrm{\scriptscriptstyle T}}\beta^{(t)}|}{\sigma^{(t)}\sqrt{c_{k}^{\mathrm{\scriptscriptstyle T}}Mc_{k}}}>\epsilon\right)$ (7)
for any given $\epsilon$. Adapting (2) to $\gamma=0$ yields $\pi(\beta,\sigma^{2}\,|\,y)=N(\beta\,|\,Mm,\sigma^{2}M)\times\text{IG}(\sigma^{2}\,|\,a,b)$, where $a=a_{0}+\frac{n}{2}$, $b=b_{0}+\frac{1}{2}\left(y^{\mathrm{\scriptscriptstyle T}}V_{y}^{-1}y+m_{0}^{\mathrm{\scriptscriptstyle T}}M_{0}m_{0}-m^{\mathrm{\scriptscriptstyle T}}Mm\right)$, $M^{-1}=M_{0}^{-1}+X^{\mathrm{\scriptscriptstyle T}}V_{y}^{-1}X$ and $m=m_{0}+X^{\mathrm{\scriptscriptstyle T}}V_{y}^{-1}y$. We draw exact posterior samples from the Normal-Gamma family. Specific domain applications may provide further insights into possible constraints on $\epsilon$ and what is considered to be a clinically meaningful difference between coefficients $\beta_{i}$ and $\beta_{j}$, which, for example, could represent mean outcomes for different treatment groups in a Bayesian Analysis Of Variance (ANOVA).
Classical p-values and the probabilities $\left\{v_{k}(\epsilon)\right\}_{k}$ defined in (6) have opposite interpretation: a small value of $p_{k}$ suggests $|c_{k}^{\mathrm{\scriptscriptstyle T}}\beta|$ is far from 00, while small $v_{k}(\epsilon)$ suggests $|c_{k}^{\mathrm{\scriptscriptstyle T}}\beta|$ is likely to be close to 00. In fact, Bayesian inference with a flat prior on $\beta$ amounts to using the classical generalized least squares estimate $\widehat{\beta}$. Hence, with a specific prior, the p-values and $\epsilon$-difference probabilities yield identical decisions on contrasts in the following sense.
Theorem 1.
Let $c_{1},\ldots,c_{K}\in\mathbb{R}^{p}$, $X\in\mathbb{R}^{n\times p}$, $y\in\mathbb{R}^{n}$, and $V_{y}\in\mathbb{R}^{n\times n}$ be fixed. Under the standard linear model $y\sim N_{n}(X\beta,\sigma^{2}V_{y})$, let $p_{k}$ be the p-value for testing the null hypothesis $H_{0}^{(k)}:c_{k}^{\mathrm{\scriptscriptstyle T}}\beta=0$ against $H_{1}^{(k)}:c_{k}^{\mathrm{\scriptscriptstyle T}}\beta\neq 0$ for $k=1,\ldots,K$. Then, $p_{k}>p_{k^{\prime}}$ if and only if $v_{k}(\epsilon)<v_{k^{\prime}}(\epsilon)$ for all $k\neq k^{\prime}$, where $v_{k}(\epsilon)=\mathbb{P}\left(\left.\frac{|c_{k}^{\mathrm{\scriptscriptstyle T}}\beta|}{\sigma\sqrt{c_{k}^{\mathrm{\scriptscriptstyle T}}Mc_{k}}}>\epsilon\;\right|\;y\right)$ under $\pi(\beta,\sigma^{2})\propto\pi(\sigma^{2})$.
Proof.
See LABEL:append:prop1_proof.
Theorem 1 implies that the ascending order of the p-values associated with null hypotheses of the form $H_{0}^{(k)}:c_{k}^{\mathrm{\scriptscriptstyle T}}\beta=0$ with two-sided alternative $H_{1}^{(k)}:c_{k}^{\mathrm{\scriptscriptstyle T}}\beta\neq 0$ is equivalent to the descending order of the Bayesian $\epsilon$-difference probabilities. Since this result holds for all $\epsilon>0$ and the classical procedure does not depend on $\epsilon$, the Bayesian rankings are stable with respect to $\epsilon$.
3.2. Extension to Spatial Random Effects
We consider (1), where each observation corresponds to one region. We let $V_{\phi}$ be known, whereas $\rho$ can either be a fixed hyperparameter or a random variable with a prior density. Generalizing the $\epsilon$-difference probabilities in the fixed effects model, we conduct posterior inference on random variables $r_{k}(\epsilon)=\text{I}\left(\frac{|c_{k}^{\mathrm{\scriptscriptstyle T}}\gamma_{\star}|}{\sigma\sqrt{c_{k}^{\mathrm{\scriptscriptstyle T}}M_{\star}c_{k}}}>\epsilon\right)$ for any given $\epsilon>0$. For $k=1\ldots,K$, we define
$h_{k}(\epsilon;\rho)=\mathbb{P}\left(\left.\frac{|c_{k}^{\mathrm{\scriptscriptstyle T}}\gamma_{\star}|}{\sigma\sqrt{c_{k}^{\mathrm{\scriptscriptstyle T}}M_{\star}c_{k}}}>\epsilon\;\right|\;y,\rho\right),$ (8)
where $c_{k}\in\mathbb{R}^{n+p}$ is a contrast on $\gamma_{\star}=(\beta^{\mathrm{\scriptscriptstyle T}},\gamma^{\mathrm{\scriptscriptstyle T}})^{\mathrm{\scriptscriptstyle T}}$. If $c_{k}^{\mathrm{\scriptscriptstyle T}}\gamma_{\star}=a_{k}^{\mathrm{\scriptscriptstyle T}}\gamma$ for some $a_{k}\in\mathbb{R}^{n}$, then $h_{k}(\epsilon;\rho)=\mathbb{P}\left(\left.\frac{|a_{k}^{\mathrm{\scriptscriptstyle T}}\phi|}{\sqrt{a_{k}^{\mathrm{\scriptscriptstyle T}}\operatorname{Var}(\phi\,|\,y,\sigma^{2},\rho)a_{k}}}>\epsilon\;\right|\;y,\rho\right)$ where $\operatorname{Var}(\phi\,|\,y,\sigma^{2},\rho)$ is given by (4).
In the context of detecting spatial disparities, consider $K$ pairs of neighboring regions $L=\{(i,j):i<j,i\sim j\}$, where $i\sim j$ means that regions $i$ and $j$ are neighbors. If $c_{ij}\in\mathbb{R}^{n+p}$ and $a_{ij}\in\mathbb{R}^{n}$ such that $c_{ij}^{\mathrm{\scriptscriptstyle T}}\gamma_{\star}=a_{ij}^{\mathrm{\scriptscriptstyle T}}\gamma=\gamma_{i}-\gamma_{j}$, then the corresponding $\epsilon$-difference probability is $h_{ij}(\epsilon;\rho)=\mathbb{P}\left(\left.\frac{|\phi_{i}-\phi_{j}|}{\sqrt{a_{ij}^{\mathrm{\scriptscriptstyle T}}\operatorname{Var}(\phi\,|\,y,\sigma^{2},\rho)a_{ij}}}>\epsilon\;\right|\;y,\rho\right)$. Using (1) and (2), we extend the insights obtained from Theorem 1 and show that the rankings of conditional posterior probabilities $h_{1}(\epsilon),\ldots,h_{K}(\epsilon)$ are stable with respect to $\epsilon$ for any $\rho\in(0,1)$.