### v4

# Assessing Spatial Disparities: A Bayesian Linear Regression Approach

Abstract.
Epidemiological investigations of regionally aggregated spatial data often involve detecting spatial health disparities among neighboring regions on a map of disease mortality or incidence rates. Analyzing such data introduces spatial dependence among health outcomes and seeks to report statistically significant spatial disparities by delineating boundaries that separate neighboring regions with disparate health outcomes. However, there are statistical challenges to appropriately define what constitutes a spatial disparity and to construct robust probabilistic inferences for spatial disparities. We enrich the familiar Bayesian linear regression framework to introduce spatial autoregression and offer model-based detection of spatial disparities. We derive exploitable analytical tractability that considerably accelerates computation. Simulation experiments conducted on a county map of the entire United States demonstrate the effectiveness of our method, and we apply our method to a data set from the Institute of Health Metrics and Evaluation (IHME) on age-standardized US county-level estimates of lung cancer mortality rates.
1. Introduction
We develop statistical methods for detecting spatial disparities on maps. Disparities in health outcomes are a rapidly evolving field within public health (see, e.g., rao2023) and spatial variation in health outcomes affects how disparities manifest on disease maps. Spatial data analysis in epidemiological investigations is extensively documented (see, e.g., waller2004applied; waller2010handbook; lawson2013statistical; lawson2016handbook, and references therein). Inference on disparities proceeds from “difference boundaries” that represent significant differences between neighbors and delineate regions with significantly different disease mortality and incidence rates, helping to improve the geographic allocation of health care resources. In geographic analysis, the problem is often called “wombling” (see jacquez2003local; jacquez2003geographic; lu2005bayesian; li2011mining; ma2007bayesian; fitzpatrick2010, for some algorithmic approaches with applications). Probabilistic model-based approaches include developments in lu2007bayesian; ma2010hierarchical; li2015bayesian; banerjee2012bayesian; hanson2015spatial; leeBoundaryDetectionDisease2012; corpas-burgos2020serra; gaoSpatialDifferenceBoundary2023 and aiello2023detecting.
The aforementioned articles present an impressive range of methods, but are broadly classified into two groups. The first embraces hierarchical models and Bayesian inference for the spatial adjacency matrix by modeling its elements. The second endows the spatial random effects with a probability law that accommodates nonzero probabilities on neighboring spatial random effects being equal (often using an areal Dirichlet process or some adaptation thereof), which are computed as posterior probabilities. Both groups introduce spatial dependence and aim to balance the underlying spatial similarities with the ability to detect differences among neighbors. Our current contribution offers a computationally scalable and analytically tractable approach to this problem. Instead of modeling spatial effects as realizations of a Dirichlet process, which struggles to learn from the data and impedes convergence of iterative estimation algorithms, we devise fully model-based Bayesian inference within an augmented spatial linear regression framework. More specifically, we treat spatial disparities as a multiple comparison problem within a Bayesian inferential paradigm by testing for differences in spatial random effects between neighboring regions. Subsequently, we use posterior probabilities of such differences to detect spatial disparities while controlling a Bayesian false discovery rate or FDR (mullerFDRBayesianMultiple2006). The benefits of our approach include clearer interpretation of spatial effects, computational simplicity, accounting for small differences in the values of neighboring spatial effects of adjacent regions in determining spatial disparity, and providing improved sensitivity and specificity for the detection of boundary differences (as seen in the analysis by gaoSpatialDifferenceBoundary2023). Our parametric approach is suitable for applications where one disease map is of interest and lack of replication hinders non-parametric methods (such as in slackGeographicConcentrationUs2014; jagai2017county; rodriguez2018county).
We exploit closed-form distribution theory to account for the differences between the values of the spatial effects and evaluate posterior probabilities of such differences exceeding a certain threshold. We extend lu2005bayesian and fitzpatrick2010, who use only posterior means or spatial residuals to detect differences, to fully model-based inference. We provide an outline for the rest of the article. Section 2 discusses the problem of detecting differences between fixed effects in a linear model before we extend to spatial random effects. Section 3 develops the notion of a difference boundary, demonstrates a meaningful connection to traditional hypothesis testing in the simple linear regression model, and considers extensions to a generalized linear model. Section LABEL:sec:simulation_experiments evaluates our methods alongside an existing Bayesian nonparametric approach using simulated data over a map of Californian counties followed by an analysis of lung cancer mortality data in Section LABEL:sec:application. Section LABEL:sec:_discussion concludes with a brief discussion.
2. Assessing Contrasts in Bayesian Linear Regression
2.1. Random Effects under an Augmented Linear System
We follow rieblerIntuitiveBayesianSpatial2016 who adapted the “BYM” model (besagBayesianImageRestoration1991) to ensure identifiable parameters and provide a better interpretation. We write this “BYM2” model as
$y=X\beta+\gamma+\eta,\;\;\gamma=\sigma\sqrt{\rho}\phi,\;\;\phi\sim N_{n}(0,V_{\phi}),\;\;\eta=\sigma\sqrt{1-\rho}v,\;\;v\sim N_{n}(0,I_{n})$ (1)
where $\phi$ is an $n\times 1$ vector of random effects, $V_{\phi}$ is a known positive definite matrix such that $\pi(\phi)$ is proper. The parameter $\sigma^{2}$ captures the overall error variance if $c$ is set so that the geometric mean of the marginal prior variances is equal to one, and $\rho$ represents the proportion of total variance attributed to the spatial structure (rieblerIntuitiveBayesianSpatial2016). For subsequent analysis, we specify a conditionally autoregressive (CAR) structure, one of the most popular spatial models used in areal data analysis. Here, $\pi(\gamma_{i}\,|\,\sigma^{2},\gamma_{j},j\neq i)=N(\alpha\sum_{j}w_{ij}\gamma_{j}/w_{i+},c\sigma^{2}/w_{i+})$, where $w_{ij}$ is the binary adjacency indicator for regions $i$ and $j$, $\alpha$ is a spatial smoothing parameter, $c$ is a fixed scaling constant and $w_{i+}=\sum_{j}w_{ij}$ is the number of neighbors of region $i$. The conditional distribution of each spatial residual is based on a linear combination of its neighbors. By Brook’s lemma (besagBayesianImageRestoration1991), this equates to setting $V^{-1}_{\phi}=c(D_{W}-\alpha W)$ in (1), where $W$ is the adjacency matrix, $D_{W}$ is diagonal with the diagonal elements $D_{W,ii}=w_{i+}$ (banerjeeHierarchicalModelingAnalysis2015, Chap. 4). The density $\pi(\phi)$ is proper when $\alpha\in(1/\lambda_{min},1)$, where $\lambda_{min}$ is the smallest eigenvalue of $D_{W}^{-1/2}WD_{W}^{-1/2}$. Since $\rho$ already controls the amount of spatial smoothing in (1), we set $\alpha=0.99$ in the following sections to improve the identifiability of spatial effects $\phi$, which we consider as potential indicators of latent factors driving local health inequity.
If $\pi(\gamma,\beta,\sigma^{2},\rho)=N(\gamma\,|\,0,\sigma^{2}\rho V_{\phi})\times N(\beta\,|\,M_{0}m_{0},\sigma^{2}M_{0})\times\text{IG}(\sigma^{2}\,|\,a_{\sigma^{2}},b_{\sigma^{2}})\times\pi(\rho)$ is the prior for (1), then the joint distribution arises from the augmented linear model,
$\underbrace{\begin{bmatrix}y\\ M_{0}m_{0}\\ 0_{n}\end{bmatrix}}_{y_{\star}}=\underbrace{\begin{bmatrix}X&I_{n}\\ I_{p}&0_{p\times n}\\ 0_{n\times p}&I_{n}\end{bmatrix}}_{X_{\star}}\underbrace{\begin{bmatrix}\beta\\ \gamma\end{bmatrix}}_{\gamma_{\star}}+\underbrace{\begin{bmatrix}\eta\\ \eta_{\beta}\\ \eta_{\phi}\end{bmatrix}}_{\eta_{\star}},\;\;\eta_{\star}\sim N_{2n+p}\left(0_{2n+p},\sigma^{2}V_{y_{\star}}\right)\;,$ (2)
where $V_{y_{\star}}=\begin{bmatrix}(1-\rho)I_{n}&0&0\\ 0&M_{0}&0\\ 0&0&\rho V_{\phi}\end{bmatrix}$ and a prior $\pi(\gamma_{\star},\sigma^{2},\rho)=\text{IG}(\sigma^{2}\,|\,a_{\sigma^{2}},b_{\sigma^{2}})\times\pi(\rho)$ (since the prior for $\gamma_{\star}=(\beta^{\mathrm{\scriptscriptstyle T}},\gamma^{\mathrm{\scriptscriptstyle T}})^{\mathrm{\scriptscriptstyle T}}$ has been absorbed into the linear system). The resulting conditional posterior joint distribution, $\pi(\gamma_{\star},\sigma^{2}\,|\,y,\rho)$, is
${}\pi(\gamma_{\star},\sigma^{2}\,|\,y,\rho)=\text{N}_{n+p}\left(\gamma_{\star}\,|\,M_{\star}m_{\star},\sigma^{2}M_{\star}\right)\times\text{IG}\left(\sigma^{2}\,|\,a_{n},b_{n}\right)\;,$ (3)
where $M_{\star}=(X_{\star}^{\mathrm{\scriptscriptstyle T}}V_{y_{\star}}^{-1}X_{\star})^{-1}$, $m_{\star}=X_{\star}^{\mathrm{\scriptscriptstyle T}}V_{y_{\star}}^{-1}y_{\star}=\left(\frac{1}{1-\rho}y^{\mathrm{\scriptscriptstyle T}}X+m_{0}^{\mathrm{\scriptscriptstyle T}},\frac{1}{1-\rho}y^{\mathrm{\scriptscriptstyle T}}\right)^{\mathrm{\scriptscriptstyle T}}$, $a_{n}=a_{\sigma^{2}}+\frac{n}{2}$, and $b_{n}=b_{\sigma^{2}}+\frac{1}{2}\left(\frac{1}{1-\rho}y^{\mathrm{\scriptscriptstyle T}}y+m_{0}^{\mathrm{\scriptscriptstyle T}}M_{0}m_{0}-m_{\star}^{\mathrm{\scriptscriptstyle T}}M_{\star}m_{\star}\right)$. The posterior mean $M_{\star}m_{\star}=\widehat{\gamma}_{\star}$ is the general least squares regression estimate to be solved $X_{\star}^{\mathrm{\scriptscriptstyle T}}V_{y_{\star}}^{-1}X_{\star}\widehat{\gamma}_{\star}=X_{\star}^{\mathrm{\scriptscriptstyle T}}V_{y_{\star}}^{-1}y_{\star}$ and $b_{n}=b_{\sigma^{2}}+\frac{ns^{2}}{2}$, where $s^{2}=\frac{1}{n}(y_{\star}-X_{\star}\widehat{\gamma}_{\star})^{\mathrm{\scriptscriptstyle T}}V_{y_{\star}}^{-1}(y_{\star}-X_{\star}\widehat{\gamma}_{\star})$ is the mean residual sums of squares from (2).
If $\rho\in(0,1)$ is fixed, then we can draw exact posterior samples from (3). Noting that $\gamma=\sigma\sqrt{\rho}\phi$ and carrying out some familiar linear algebra calculations, we obtain
$\operatorname{Var}(\phi\,|\,y,\sigma^{2},\rho)=\left(V_{\phi}^{-1}+\frac{\rho}{1-\rho}\left(I_{n}-XA^{-1}X^{\mathrm{\scriptscriptstyle T}}\right)\right)^{-1},$ (4)
where $A=X^{\mathrm{\scriptscriptstyle T}}X+(1-\rho)M_{0}^{-1}$. A flat prior to $\beta$, that is, $M_{0}^{-1}=O$, yields $A=X^{\mathrm{\scriptscriptstyle T}}X$ and
$\lim_{M_{0}^{-1}\to O}\operatorname{Var}(\phi\,|\,y,\sigma^{2},\rho)=\left(V_{\phi}^{-1}+\frac{\rho}{1-\rho}\left(I_{n}-H\right)\right)^{-1},\;\mbox{ where }\;H=X(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}\;.$ (5)
These expressions are used in Section 3.2 to formulate the detection of differences between boundaries.
2.2. Limiting cases
The limits of the conditional posterior distribution of $(\beta,\gamma)$ as $\rho\rightarrow 0^{+}$ or as $\rho\rightarrow 1^{-}$ are analytically derived when the prior is non-informative on $\beta$. By spectral decomposition, $V^{1/2}_{\phi}(I-H)V^{1/2}_{\phi}=CDC^{\mathrm{\scriptscriptstyle T}}$ where $C$ is an orthogonal matrix and $D=\text{diag}(d_{1},\ldots,d_{n})$ with entries that are nonnegative eigenvalues of $V_{\phi}^{1/2}\left(I_{n}-H\right)V_{\phi}^{1/2}$. Let $U=V^{1/2}_{\phi}C$, then $U$ is invertible, $U^{\mathrm{\scriptscriptstyle T}}V^{-1}_{\phi}U=I_{n}$, $U^{\mathrm{\scriptscriptstyle T}}(I_{n}-H)U=D$, $\operatorname{Var}(\gamma\,|\,y,\sigma^{2},\rho)=\sigma^{2}U\left(\frac{1}{\rho}I_{n}+\frac{1}{1-\rho}D\right)^{-1}U^{\mathrm{\scriptscriptstyle T}}$, and $\mathbb{E}[\gamma\,|\,y,\sigma^{2},\rho]=U\left(\frac{1-\rho}{\rho}I_{n}+D\right)^{-1}DU^{-1}y$. Calculations reveal $\lim_{\rho\rightarrow 0^{+}}\operatorname{Var}(\gamma\,|\,y,\sigma^{2},\rho)=0_{n\times n}$ and $\lim_{\rho\rightarrow 0^{+}}\mathbb{E}[\gamma\,|\,y,\sigma^{2},\rho]=0_{n}$. Therefore, if the prior on $\beta$ is flat, then $\lim_{\rho\rightarrow 0^{+}}\pi(\beta,\gamma,\sigma^{2}\,|\,y,\rho)$ collapses to the degenerate mass at $\gamma=0_{n}$.
We also obtain $\lim_{\rho\rightarrow 1^{-}}\left(\frac{1}{\rho}I_{n}+\frac{1}{1-\rho}D\right)^{-1}=I_{n}-D^{\star}$ and $\lim_{\rho\rightarrow 1^{-}}\left(\frac{1-\rho}{\rho}I_{n}+D\right)^{-1}D=D^{\star}$ where $D^{\star}\in\mathbb{R}^{n\times n}$ is diagonal with the entries $D_{ii}^{\star}=I(d_{i}>0)$ for $i=1,\ldots,n$. Further simplifications reveal that $\mathbb{E}[\beta\,|\,y,\sigma^{2},\rho]\rightarrow(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}\left(y-UD^{\star}U^{-1}e\right)$ and $\mathbb{E}[\gamma\,|\,y,\sigma^{2},\rho]\rightarrow UD^{\star}U^{-1}e$ as $\rho\rightarrow 1^{-}$, where $e=(I-H)y$ are the residuals of ordinary least squares. Furthermore, $\operatorname{Var}(\beta\,|\,y,\sigma^{2},\rho)\rightarrow\sigma^{2}(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}VX(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}$, $\operatorname{Var}(\gamma\,|\,y,\sigma^{2},\rho)\rightarrow\sigma^{2}V$ and $\operatorname{Cov}(\beta,\gamma\,|\,y,\sigma^{2},\rho)\rightarrow-\sigma^{2}(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}V$ as $\rho\rightarrow 1^{-}$, where $V=V_{\phi}-UD^{\star}U^{\mathrm{\scriptscriptstyle T}}$. Finally, $\lim_{\rho\rightarrow 1^{-}}b_{n}=b_{\sigma^{2}}+\frac{1}{2}v^{\mathrm{\scriptscriptstyle T}}D^{\star}v$, where $v=U^{-1}e$. Therefore, $\lim_{\rho\rightarrow 1^{-}}\pi(\sigma^{2}\,|\,y,\rho)=\text{IG}(\sigma^{2}\,|\,a_{\sigma^{2}}+n/2,b_{\sigma^{2}}+v^{\mathrm{\scriptscriptstyle T}}D^{\star}v/2)$. These limiting cases have implications for subsequent boundary detection (see Section LABEL:subsec:epsilon_intro).
3. A New Framework for Detecting Spatial Disparities
3.1. Connection to Classical Hypothesis Testing
We first briefly address contrast testing in Bayesian linear regression without spatial effects, that is, $\gamma=0$ in (1). Whereas $\mathbb{P}(c_{k}^{\mathrm{\scriptscriptstyle T}}\beta=0\,|\,y)=0$ for a continuous posterior distribution, we can test $H^{(k)}_{0}:c_{k}^{\mathrm{\scriptscriptstyle T}}\beta=0$ against $H^{(k)}_{1}:c_{k}^{\mathrm{\scriptscriptstyle T}}\beta\neq 0$ for $K$ nonzero $p\times 1$ vectors $c_{1},\ldots,c_{K}$ using a classical linear model, $y=X\beta+\eta,\eta\sim N_{n}(0_{n},\sigma^{2}V_{y})$, where $\beta$ is fixed but unknown (e.g., seberHypothesisTesting2003). We compute $t_{\text{obs}}^{(k)}=\frac{c_{k}^{\mathrm{\scriptscriptstyle T}}\widehat{\beta}}{\widehat{\sigma}\sqrt{c_{k}^{\mathrm{\scriptscriptstyle T}}(X^{\mathrm{\scriptscriptstyle T}}V_{y}^{-1}X)^{-1}c_{k}}}$ for each $c_{k}$, where $\widehat{\beta}=(X^{\mathrm{\scriptscriptstyle T}}V_{y}^{-1}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}V_{y}^{-1}y$ and $(n-p)\widehat{\sigma}^{2}=\left(y-X\widehat{\beta}\right)^{\mathrm{\scriptscriptstyle T}}V_{y}^{-1}\left(y-X\widehat{\beta}\right)$ are generalized least squares estimates. We reject $H^{(k)}_{0}$ if the p-values $p_{k}=\mathbb{P}\left(T>\left|t^{(k)}_{\text{obs}}\right|\right)$, where $T\sim t_{n-p}$, are below a multiplicity adjusted threshold.
For Bayesian inference, we retain the continuous conjugate prior on $\beta$ and evaluate the posterior probabilities for the absolute standardized linear combinations between elements of $\beta$ to exceed a threshold $\epsilon$. Hence, for all $k=1,\ldots,K$ and any $\epsilon>0$, we define
$v_{k}(\epsilon)=\mathbb{P}\left(\left.\frac{|c_{k}^{\mathrm{\scriptscriptstyle T}}\beta|}{\sigma\sqrt{c_{k}^{\mathrm{\scriptscriptstyle T}}Mc_{k}}}>\epsilon\;\right|\;y\right).$ (6)
We refer to these posterior probabilities as $\epsilon$-difference probabilities, where $\epsilon$ represents a universal threshold across all $K$ linear combinations indicating the number of posterior standard deviation units that $|c_{k}^{\mathrm{\scriptscriptstyle T}}\beta|$ is away from 0. If $\beta^{(1)},\ldots,\beta^{(N)}$ and $\sigma^{(1)},\ldots,\sigma^{(N)}$ are $N$ samples drawn from their respective posterior distributions, then we estimate (6) using
$\widehat{v}_{k}(\epsilon)=\frac{1}{N}\sum_{t=1}^{N}\text{I}\left(\frac{|c_{k}^{\mathrm{\scriptscriptstyle T}}\beta^{(t)}|}{\sigma^{(t)}\sqrt{c_{k}^{\mathrm{\scriptscriptstyle T}}Mc_{k}}}>\epsilon\right)$ (7)
for any given $\epsilon$. Adapting (2) to $\gamma=0$ yields $\pi(\beta,\sigma^{2}\,|\,y)=N(\beta\,|\,Mm,\sigma^{2}M)\times\text{IG}(\sigma^{2}\,|\,a,b)$, where $a=a_{0}+\frac{n}{2}$, $b=b_{0}+\frac{1}{2}\left(y^{\mathrm{\scriptscriptstyle T}}V_{y}^{-1}y+m_{0}^{\mathrm{\scriptscriptstyle T}}M_{0}m_{0}-m^{\mathrm{\scriptscriptstyle T}}Mm\right)$, $M^{-1}=M_{0}^{-1}+X^{\mathrm{\scriptscriptstyle T}}V_{y}^{-1}X$ and $m=m_{0}+X^{\mathrm{\scriptscriptstyle T}}V_{y}^{-1}y$. We draw exact posterior samples from the Normal-Gamma family. Specific domain applications may provide further information on possible constraints on $\epsilon$ and what is considered to be a clinically significant difference between coefficients $\beta_{i}$ and $\beta_{j}$, which, for example, could represent mean results for different treatment groups in a Bayesian Analysis Of Variance (ANOVA).
The classical p-values and the probabilities $\left\{v_{k}(\epsilon)\right\}_{k}$ defined in (6) have opposite interpretation: a small value of $p_{k}$ suggests that $|c_{k}^{\mathrm{\scriptscriptstyle T}}\beta|$ is far from 0, while small $v_{k}(\epsilon)$ suggests that $|c_{k}^{\mathrm{\scriptscriptstyle T}}\beta|$ is likely to be close to 0. In fact, Bayesian inference with a flat prior on $\beta$ amounts to using the classical generalized least squares estimate $\widehat{\beta}$. Hence, with a specific prior, the p-values and $\epsilon$-difference probabilities yield identical decisions on contrasts in the following sense.
Theorem 1.
Let $c_{1},\ldots,c_{K}\in\mathbb{R}^{p}$, $X\in\mathbb{R}^{n\times p}$, $y\in\mathbb{R}^{n}$ and $V_{y}\in\mathbb{R}^{n\times n}$ be fixed. Under the standard linear model $y\sim N_{n}(X\beta,\sigma^{2}V_{y})$, let $p_{k}$ be the p-value to test the null hypothesis $H_{0}^{(k)}:c_{k}^{\mathrm{\scriptscriptstyle T}}\beta=0$ against $H_{1}^{(k)}:c_{k}^{\mathrm{\scriptscriptstyle T}}\beta\neq 0$ for $k=1,\ldots,K$. Then, $p_{k}>p_{k^{\prime}}$ if and only if $v_{k}(\epsilon)<v_{k^{\prime}}(\epsilon)$ for all $k\neq k^{\prime}$, where $v_{k}(\epsilon)=\mathbb{P}\left(\left.\frac{|c_{k}^{\mathrm{\scriptscriptstyle T}}\beta|}{\sigma\sqrt{c_{k}^{\mathrm{\scriptscriptstyle T}}Mc_{k}}}>\epsilon\;\right|\;y\right)$ under $\pi(\beta,\sigma^{2})\propto\pi(\sigma^{2})$.
Proof.
See Appendix LABEL:append:prop1_proof.
Theorem 1 implies that the ascending order of the p-values associated with null hypotheses of the form $H_{0}^{(k)}:c_{k}^{\mathrm{\scriptscriptstyle T}}\beta=0$ with the two-sided alternative $H_{1}^{(k)}:c_{k}^{\mathrm{\scriptscriptstyle T}}\beta\neq 0$ is equivalent to the descending order of the Bayesian $\epsilon$-difference probabilities. Since this result holds for all $\epsilon>0$ and the classical procedure does not depend on $\epsilon$, the Bayesian rankings are stable with respect to $\epsilon$.
3.2. Extension to Spatial Random Effects
We consider (1), where each observation corresponds to one region. We let $V_{\phi}$ be known, while $\rho$ can be either a fixed hyperparameter or a random variable with a prior density. Generalizing the $\epsilon$-difference probabilities in the fixed effects model, we perform posterior inference on random variables $r_{k}(\epsilon)=\text{I}\left(\frac{|c_{k}^{\mathrm{\scriptscriptstyle T}}\gamma_{\star}|}{\sigma\sqrt{c_{k}^{\mathrm{\scriptscriptstyle T}}M_{\star}c_{k}}}>\epsilon\right)$ for any given $\epsilon>0$. For $k=1\ldots,K$, we define
$h_{k}(\epsilon;\rho)=\mathbb{P}\left(\left.\frac{|c_{k}^{\mathrm{\scriptscriptstyle T}}\gamma_{\star}|}{\sigma\sqrt{c_{k}^{\mathrm{\scriptscriptstyle T}}M_{\star}c_{k}}}>\epsilon\;\right|\;y,\rho\right),$ (8)
where $c_{k}\in\mathbb{R}^{n+p}$ is a contrast on $\gamma_{\star}=(\beta^{\mathrm{\scriptscriptstyle T}},\gamma^{\mathrm{\scriptscriptstyle T}})^{\mathrm{\scriptscriptstyle T}}$. If $c_{k}^{\mathrm{\scriptscriptstyle T}}\gamma_{\star}=a_{k}^{\mathrm{\scriptscriptstyle T}}\gamma$ for some $a_{k}\in\mathbb{R}^{n}$, then $h_{k}(\epsilon;\rho)=\mathbb{P}\left(\left.\frac{|a_{k}^{\mathrm{\scriptscriptstyle T}}\phi|}{\sqrt{a_{k}^{\mathrm{\scriptscriptstyle T}}\operatorname{Var}(\phi\,|\,y,\sigma^{2},\rho)a_{k}}}>\epsilon\;\right|\;y,\rho\right)$ where $\operatorname{Var}(\phi\,|\,y,\sigma^{2},\rho)$ is given by (4).
In the context of detecting spatial disparities, consider $K$ pairs of neighboring regions $L=\{(i,j):i<j,i\sim j\}$, where $i\sim j$ means that the regions $i$ and $j$ are neighbors. If $c_{ij}\in\mathbb{R}^{n+p}$ and $a_{ij}\in\mathbb{R}^{n}$ are such that $c_{ij}^{\mathrm{\scriptscriptstyle T}}\gamma_{\star}=a_{ij}^{\mathrm{\scriptscriptstyle T}}\gamma=\gamma_{i}-\gamma_{j}$, then the corresponding $\epsilon$-difference probability is $h_{ij}(\epsilon;\rho)=\mathbb{P}\left(\left.\frac{|\phi_{i}-\phi_{j}|}{\sqrt{a_{ij}^{\mathrm{\scriptscriptstyle T}}\operatorname{Var}(\phi\,|\,y,\sigma^{2},\rho)a_{ij}}}>\epsilon\;\right|\;y,\rho\right)$. Using (1) and (2), we extend the insights obtained from Theorem 1 and show that the ranking of conditional posterior probabilities $h_{1}(\epsilon),\ldots,h_{K}(\epsilon)$ is stable with respect to $\epsilon$ for any $\rho\in(0,1)$.